# Formation of clusters and coarsening in weakly interacting diffusions

### v1

@@ 3.4. Multi-cluster states
When after initial clustering, the $N$ particles in (1.1) are divided into $K\in\mathbb{N}$ clusters of $N^{(1)},\dots,N^{(K)}$ particles, respectively, we can assign to each cluster a mass $m_{i}=\frac{N^{(i)}}{N}$. If the clusters are sufficiently localized, the particles within each cluster behave approximately like the system (1.1) with effective interaction strengths $\gamma_{i}=\frac{N^{(i)}}{N}\gamma=m^{(i)}\gamma$. Therefore, the particles within cluster $i$ are approximately normally distributed with variances $\sigma_{i}^{2}=\frac{\ell}{\gamma w^{\prime\prime}(0)m^{(i)}}$. Moreover, the empirical density for a multi-cluster state with masses $m^{(1)},\dots,m^{(K)}$ and cluster centers $X^{(1)},\dots,X^{(K)}$ is approximately a Gaussian mixture
@@ 3.5. The critical minimal mass for a cluster inside a multi-cluster state
Consider again a multi-cluster state in the particle system (1.1). In order for a cluster of mass $m\in(0,1)$ of this multi-cluster state to be stable, the width $\sigma_{\rm cluster}$ needs to be small compared to the interaction range $2s_{w}\ell$ of the potential since otherwise the cluster would dissolve quickly. This gives the condition
$m>m_{\rm crit}\asymp\frac{1}{\gamma\ell w^{\prime\prime}(0)}$ (3.41)
for a cluster of mass $m$ to be stable, where $m_{\rm crit}$ is the critical minimal mass. This can be also seen through the following scaling argument for the free energy:
@@
Therefore, $\rho_{m,\gamma}$ can only be a stable cluster state provided that $m\gamma>\gamma_{c}$, where $\gamma_{c}$ is defined by Section 3.1. This suggests that the critical mass $m_{\rm crit}\in(0,1)$ is of the form $m_{\rm crit}=\frac{\gamma_{c}}{\gamma}$, so that (3.41) gives
$\gamma_{c}\asymp\frac{1}{\ell w^{\prime\prime}(0)}.$ (3.46)
@@ 4.1. The Markovian dynamics of the clusters in the particle model after [GPY17]
We now describe the coalescense of the clusters for the particle model (1.1), following closely [GPY17]. Assuming that $w$ satisfies 1.1, we consider again the dynamics (1.1) in the regime $\ell\ll 1$ and $\gamma\gg 1$. Let us assume that initial clustering has already occured (see Section 2) so that the particles are organized in $K(\tau_{0})$ clusters at some time $\tau_{0}$. Let $I^{j}\subset\{1,\dots,N\}$ be the set of indices of the particles belonging to the $j$-th cluster. Each cluster $j$ has has then a mass of $m^{(j)}=\lvert I^{j}\rvert/N$ and a corresponding width $\sigma_{j}$ defined by $\sigma^{2}_{j}=\frac{\ell}{\gamma w^{\prime\prime}(0)m^{(j)}}$. Since $w$ is even, it follows from (1.1) that the cluster centers
@@
where $W^{(j)}_{t}:=\frac{1}{\sqrt{\lvert I^{j}\rvert}}\sum_{i\in I^{j}}W^{i}_{t}$ is a standard Brownian motion. The cluster centers follow (4.2) until two clusters meet at time
@@
where the characteristic length scale $d$ is given by $d\asymp\ell s_{w}$. In this case, the two clusters $k$ and $l$ merge expect for some set of realizations which is exponentially small in $\gamma$. It follows from (1.1) that the merging process can be described approximately by the following SDE
${\rm d}big(X^{(j)}_{t}big)=-\frac{\gamma w^{\prime\prime}(0)}{2}big(m^{(j)}+m^{(k)}big)big(X^{(j)}_{t}-X^{(k)}_{t}big)\mathop{}\!\mathup{d}t$ (4.4)
and similarly for $X^{(k)}_{t}$. In particular, the cluster centers converge exponentially fast. The new cluster with index set $I^{j}\cup I^{k}$ has mass
@@
where $W^{(j,k)}_{t}:=\frac{1}{\sqrt{2}}\left(W^{(j)}_{t}+W^{(k)}_{t}\right)$ is a standard Brownian motion and the center of the new cluster starts at some time $\tau_{1}$ at the weighted average
@@
Relabeling the clusters to take the merging into account, we obtain now $K(\tau_{1})=K(\tau_{0})-1$ clusters. The coalescing process continues until a unique cluster with mass $1$ and variance $\frac{\ell}{\gamma w^{\prime\prime}(0)}$ forms. Neglecting the set of realizations on which merging of touching clusters does not occur, or mass leakage occurs, it follows from standard properties [^2] For a standard Brownian motion $W_{t}$, the hitting time $\tau_{a,-b}$ of $\{a,-b\}$ (with $a,b>0$) satisfies $\mathbf{E}\left[\tau_{a,-b}\right]=ab$. Therefore, from (4.2) we see that for two clusters on a periodic domain $[0,1]$, the timescale for two clusters to merge is of order $N$. of the one dimensional Brownian motion in a bounded domain that a unique cluster will form almost surely in finite time on a timescale of order $t_{\rm coalesence}\asymp N$.
@@ 4.2. The extended Markovian dynamics with mass exchange
We now extend the model of the Markovian dynamics of the clusters by allowing the clusters to exchange mass. Under 1.1, we propose the following model, which we derive in Section 4.5: We assume that at time $t$, there are $K(t)$ clusters with centers $X^{(1)}<\dots<X^{(K(t))}$ and masses $m^{(1)},\dots,m^{(K(t))}\in(0,1)$ such that $\sum_{j=1}^{K(t)}m^{(j)}=1$. The positions and the masses of the clusters behave as follows:
@@
  • Cluster merging: If two clusters with centers $X^{(j)}$ and $X^{(k)}$ meet, they merge into a single cluster of mass $m^{(j)}+m^{(k)}$ and position given by (4.7), so that the system (4.8) starts anew with the clusters $X^{(j)}$ and $X^{(k)}$ being replaced by a single cluster with starting position $X^{(j,k)}$ and mass $m^{(j)}+m^{(k)}$.
@@ 4.3. Application to the mean-field equation and dynamic metastability
The behaviour of the mean-field PDE (1.6) after initial clustering in the regime $\gamma\ell\gg 1$ and $\ell\ll 1$ is governed by the limit $N\to\infty$ of (4.8), so that the cluster centers are frozen in space and the mass exchange ODE system (4.8b) determines the dynamics. This is also reflected by the coalescence timescale $t_{\rm coalesence}\asymp N$. This suggests that after initial clustering, coarsening in the PDE (1.6) occurs by mass exchange between clusters which is described by (4.8b). Indeed, if we start the PDE with a multi-clustered initial state, then the ODE (4.8b) approximately describes the PDE evolution, as shown in Fig. 4.1 for the potential (1.4). Fig. 4.2 shows a similar comparison for the truncated piecewise parabolic potential with parameters $\alpha,\beta>0$, $a\in(0,1)$ given by
@@
For $\gamma\ell\gg 1$, the timescales $e^{\gamma\ell\Delta m}$ (see (LABEL:eq:clusters-masses:Eyring-Kramers) below) for a particle to leave a cluster are exponentially large in the cluster masses $m^{(j)}$, so that we expect dynamical metastability for the cluster mass dynamics. Indeed, assume for instance that we start the PDE with two clusters with masses $m_{a},m_{b}\approx\frac{1}{2}$ and $m_{a}>m_{b}$. In the regime $\gamma\ell\gg 1$, $\ell\ll 1$, we find from (4.8b) that mass is exchanged at a rate that is exponentially small in $\gamma\ell m_{a}$ and $\gamma\ell m_{b}$. Therefore, it will take exponentially long for the stronger cluster to absorb the smaller one. Indeed, we observe both for the PDE and the mass exchange ODE (4.8b) that the masses change only very little over a large amount of time to then suddenly collapse to a state in which only the initially heavier cluster remains, see Fig. 4.1.
@@
Observe that the model is only a valid approximate description of the PDE with a multi-clustered initial state if $X^{(j+1)}-X^{(j)}\gg\ell s_{w}$, that is, the initial centers of the clusters are sufficiently far apart, since otherwise deterministic merging phenomena.
Moreover, the model (4.8b) is in general not a good approximation of the dissolution of clusters. Once one of the clusters has a mass smaller than the critical mass $m_{\rm crit}$ (see Section 3.5), then the dynamics of the particles of this cluster are not described by the Ornstein–Uhlenbeck approximation from [GPY17] and the cluster shape is in general not well approximated by (3.40). Moreover, in this situation, not only the behavior of $w$ around zero but its general shape determines the behavior and the exact evolution of mass. In particular, the description of the ODE (4.8b) is in general only valid until the first dissolution of a cluster.
@@
As before, we assume that for some fixed time $t\geq 0$, the particles (1.1) are arranged in clusters of masses $m^{(1)},\dots,m^{(K(t))}\in(0,1)$ and centers $X^{(1)}<\dots<X^{(K(t))}$. For simplicity, let us assume that the particles move on the real line $\mathbb{R}$ instead of the torus. Each cluster $X^{(j)}$ consists of approximately $m^{(j)}N$ particles. We would like to compute the timescale of a particle $X^{i}_{t}$ in the $j$-th cluster to leave the cluster and join another one, where $i\in\{1,\dots,N\}$ and $j\in\{1,\dots,K(t)\}$. The empirical measure $\mu^{N}_{t}:=\frac{1}{N}\sum_{i=1}^{N}\delta_{X^{i}_{t}}$ can be approximated by a linear combination of Gaussian densities centered at the cluster centers $X^{(k)}_{t}$ with variances $\sigma^{2}_{k}=\frac{\ell}{\gamma w^{\prime\prime}(0)m^{(k)}}$ and weights $m^{(k)}$:
@@
where $g(x;\mu,\sigma^{2})$ is the density of a normal distribution with mean $\mu$ and variance $\sigma^{2}$. We are interested in the event that the particle $X^{i}_{t}$ of cluster $j$ leaves this cluster and joins another cluster which can be assumed to be one of the neighboring clusters j1j-1 or $j+1$, since the event that a particle jumps to a cluster that is not a direct neighbor of the cluster where the particle have a probability of lower order. The particle $X^{i}_{t}$ moves in the effective potential
$V_{\rm eff}(x):=\left(W_{\gamma,\ell}\ast\mu^{N}_{t}\right)(x)\approx\sum_{k=1}^{K(t)}m^{(k)}\left(\gamma\ell w\left(\frac{{\color[rgb]{.75,.75,.75}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.75,.75}\pgfsys@color@gray@stroke{.75}\pgfsys@color@gray@fill{.75}{\bullet}}}{\ell}\right)\ast g({\color[rgb]{.75,.75,.75}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.75,.75}\pgfsys@color@gray@stroke{.75}\pgfsys@color@gray@fill{.75}{\bullet}};X^{(k)},\sigma_{k}^{2})\right)(x),$ (4.15)
that is ${\rm d}X^{i}_{t}=-V^{\prime}_{\rm eff}(X^{i}_{t}){\rm d}t+\sqrt{2}{\rm d}B^{i}_{t}$.
In order to approximate $V_{\rm eff}$ around its local minimum $X^{(j)}$, we replace $w(x)$ in (4.15) by the quadratic approximation $w(x)\approx-\Delta+w^{\prime\prime}(0)\frac{x^{2}}{2}$ for small $\lvert x\rvert$, where $\Delta:=-\inf w=-w(0)$. This is justified by the fact that the $\sigma_{k}^{2}$ is of order $\ell/\gamma$, while the interaction range of the potential $W_{\gamma,\ell}$ is of order $\ell\gg\ell/\gamma$. Using this approximation, we obtain then
@@
for $x$ close to $X^{(k)}$. Fix some $a\in\mathbb{R}$ inside the potential well of $V_{\rm eff}$ around $X^{(j-1)}$ which is close enough to the right boundary of this potential well and some $b\in\mathbb{R}$ inside the well of $V_{\rm eff}$ around $X^{(j+1)}$ which is close enough to the left boundary of this well. Let $\tau_{a,b}$ be the time of $X^{i}_{t}$ to hit $(-\infty,a]\cup[b,\infty)$ and set $u(x)=\mathbb{E}\left[\tau_{a,b}\nonscript\>\middle|\nonscript\>\mathopen{}X^{i}_{0}=x\right]$. By classical results [DZ01, Theorem 5.7.3] [Pav14, Sec. 7.3], $u$ satisfies the boundary value problem
@@
Figure 4.3. The effective potential $V_{\rm eff}$ for the Hegselmann–Krause prototype (1.4)
We would like to apply Laplace’s method for the limit $\gamma\to\infty$. Around its minima $X^{(1)},\dots,X^{(K(t))}$, we approximate $V_{\rm eff}$ by (4.5). Also relevant is the width of the potential wells around the minima. Since the support of $W_{\gamma,\ell}$ has length $2\ell s_{w}$, we expect that the width of the potential wells of $V_{\rm eff}=W_{\gamma,\ell}\ast\mu^{N}_{t}$ can roughly be approximated by $2\ell s_{w}$, since $\sigma_{k}=\sqrt{\ell/(\gamma w^{\prime\prime}(0)m^{(k)})}\ll\ell$. For fixed $x\in(X^{(j)}-\ell s_{w},X^{(j)}+\ell s_{w})$, we therefore obtain the rough approximation
$\int_{a}^{x}e^{V_{\rm eff}(y)}\mathop{}\!\mathup{d}y$ $\approx X^{(j)}-X^{(j-1)}-2s_{w}\ell,$ (4.19)
$\int_{a}^{b}e^{V_{\rm eff}(y)}\mathop{}\!\mathup{d}y$ $\approx X^{(j+1)}-X^{(j-1)}-4s_{w}\ell.$ (4.20)
Furthermore, from (4.5) we have for $c\in(X^{(j)}-\ell s_{w},X^{(j)}+\ell s_{w})$ that
$\int_{a}^{x}e^{-V_{\rm eff}(z)}\mathop{}\!\mathup{d}z\sim\sqrt{\frac{2\pi\sigma_{j}^{2}}{e}}e^{e\gamma\ell\Delta m^{(j)}}=\sqrt{\frac{2\pi\ell}{e\gamma w^{\prime\prime}(0)m^{(j)}}}e^{\gamma\ell\Delta m^{(j)}}.$ (4.21)
A similar argument shows for the same $x$ the rough approximation
$\int_{a}^{b}\int_{a}^{y}e^{V_{\rm eff}(y)-V_{\rm eff}(z)}\mathop{}\!\mathup{d}z\mathop{}\!\mathup{d}y$ $\approx\left(X^{(j+1)}-X^{(j)}-2s_{w}\ell\right)\sqrt{\frac{2\pi\ell}{e\gamma w^{\prime\prime}(0)m^{(j)}}}e^{\gamma\ell\Delta m^{(j)}}.$ (4.22)
Here we used that $a,b$ are sufficiently close to the boundary of their respective valleys, so that we get only the contribution of the valley around $X^{(j)}$. Therefore, we obtain the Eyring–Kramers formula
  E[ τ_a,b  — X^i_0 = x]
≈ (X(j+1)- X(j)-2swℓ) ⋅(X(j)- X(j-1)-2swℓ) (X(j+1)- X(j-1)- 2swℓ) 2πℓe γw”(0) m(j) e^γℓΔm^(j)
as $\gamma\to\infty$ for all $x\in(X^{(j)}-\ell,X^{(j)}+\ell)$. Here, we used that the first term in (4.18) is of lower order than the second by our choice of $a$ and $b$. The formula (LABEL:eq:clusters-masses:Eyring-Kramers) holds for all $j=1,\dots,K(t)$, where we adopt the notation $X^{(0)}=X^{(K(t))}$ and $X^{(K(t)+1)}=X^{(1)}$ as before.
The equilibrium potential
Let $h_{a,b}(x):=\mathbb{P}\left[\tau_{a}<\tau_{b}\nonscript\>\middle|\nonscript\>\mathopen{}X^{i}_{0}=x\right]$ be the probability that a particle $X^{i}_{t}$ which starts at $x$ hits $a$ before hitting $b$. Then, $h_{a,b}$ satisfies the boundary value problem given by (see e.g. [Ber13])

### v2

@@ 3.5. Multi-cluster states
When, after initial clustering, the $N$ particles in (1.1) are divided into $K\in\mathbb{N}$ clusters of $N^{(1)},\dots,N^{(K)}$ particles, respectively, we can assign to each cluster a mass $m_{i}=\frac{N^{(i)}}{N}$. If the clusters are sufficiently localised, the particles within each cluster behave approximately like the system (1.1) with effective interaction strengths $\gamma_{i}=\frac{N^{(i)}}{N}\gamma=m^{(i)}\gamma$. Therefore, the particles within cluster $i$ are approximately normally distributed with variances $\sigma_{i}^{2}=\frac{\ell}{\gamma w^{\prime\prime}(0)m^{(i)}}$. Moreover, the empirical density for a multi-cluster state with masses $m^{(1)},\dots,m^{(K)}$ and cluster centres $X^{(1)},\dots,X^{(K)}$ is approximately a Gaussian mixture
@@ 3.6. The critical minimal mass for a cluster inside a multi-cluster state
Consider again a multi-cluster state in the particle system (1.1). In order for a cluster of mass $m\in(0,1)$ of this multi-cluster state to be stable at least for some time, the width $\sigma_{\rm cluster}=(\ell/(\gamma w^{\prime\prime}(0)m))^{\frac{1}{2}}$ needs to be small compared to the interaction range $2s_{w}\ell$ of the potential since otherwise the cluster would dissolve quickly. We define the critical minimal mass $m_{\rm crit}\in(0,1)$ as the smallest mass for which a cluster within a multi-cluster state does not dissolve immediately. Certainly, if the cluster width $\sigma_{\rm cluster}$ is much smaller than the interaction range $2s_{w}\ell$ of $W_{\gamma,\ell}$, we expect the cluster to be stable for some time. This corresponds to the condition
$\sigma_{\rm cluster}\ll 2s_{w}\ell\quad\Rightarrow\quad m\gg\frac{1}{\gamma\ell},$ (3.62)
giving the heuristic lower bound
$m_{\rm crit}\gtrsim\frac{1}{\gamma\ell}$ (3.63)
on the critical minimal mass. The existence of a critical minimal mass $m_{\rm crit}$ can also be seen through the following scaling argument for the free energy. For all $m\in(0,1)$ and $\rho\in{\mathcal{P}}(\mathbb{T})$, we have that
@@
Therefore, $\rho_{m,\gamma}$ can only be a stable cluster state provided that $m\gamma>\gamma_{c}$, where $\gamma_{c}$ is defined in Section 3.1. This suggests that the critical mass $m_{\rm crit}\in(0,1)$ is of the form $m_{\rm crit}=\frac{\gamma_{c}}{\gamma}$, so that (3.63) gives the heuristic lower bound $\gamma_{c}\gtrsim\ell^{-1}$. Recall that Proposition 3.10 gives a rigorous upper bound on $\gamma_{c}$ which is of the order $\ell^{-1}\log(\ell^{-1})$.
@@ 4.1. The Markovian dynamics of the clusters in the particle model after [GPY17]
We now describe the coalescence of the clusters for the particle model (1.1), following closely [GPY17]. Assuming that $w$ satisfies 1.1, we consider again the dynamics (1.1) in the regime $\ell\ll 1$ and $\gamma\gg 1$. Let us assume that initial clustering has already occurred (see Section 2) so that the particles are organised in $K(\tau_{0})$ clusters at some time $\tau_{0}$. Let $I^{j}\subset\{1,\dots,N\}$ be the set of indices of the particles belonging to the $j$-th cluster. Each cluster $j$ has has then a mass of $m^{(j)}=\lvert I^{j}\rvert/N$ and a corresponding width $\sigma_{j}$ defined by $\sigma^{2}_{j}=\frac{\ell}{\gamma w^{\prime\prime}(0)m^{(j)}}$. Since $w$ is even, it follows from (1.1) that the cluster centres
@@
where $W^{(j)}_{t}:=\frac{1}{\sqrt{\lvert I^{j}\rvert}}\sum_{i\in I^{j}}W^{i}_{t}$ is a standard Brownian motion. The cluster centres follow (4.2) until two clusters meet at time
@@
where the characteristic length scale $d$ is given by $d\asymp\ell s_{w}$. In this case, the two clusters $k$ and $l$ merge except for some set of realizations which is exponentially small in $\gamma$. It follows from (1.1) that the merging process can be described approximately by the following SDE
[…]
In particular, since the martingale term has quadratic variation of order $\mathcal{O}(t/N)$, for $N$ large enough, the drift term dominates and the cluster centres converge exponentially fast. The new cluster with index set $I^{j}\cup I^{k}$ has mass
@@
where $W^{(j,k)}_{t}:=\frac{1}{\sqrt{2}}\left(W^{(j)}_{t}+W^{(k)}_{t}\right)$ is a standard Brownian motion and the centre of the new cluster starts at some time $\tau_{1}$ at the weighted average
@@
Relabeling the clusters to take the merging into account, we obtain now $K(\tau_{1})=K(\tau_{0})-1$ clusters. The coalescing process continues until a unique cluster with mass $1$ and variance $\frac{\ell}{\gamma w^{\prime\prime}(0)}$ forms. Neglecting the set of realizations on which merging of touching clusters does not occur, or mass leakage occurs, it follows from standard properties [^2] For a standard Brownian motion $W_{t}$ and some $\sigma>0$, the hitting time $\tau_{a,-b}:=\inf\{t>0\nonscript\>|\nonscript\>\mathopen{}\sigma W_{t}\in\{a,-b\}\}$ (with $a,b>0$) satisfies $\mathbf{E}\left[\tau_{a,-b}\right]=ab/\sigma^{2}$. Therefore, since by (4.2) the difference of two cluster centres behaves (before merging) as $\sigma B_{t}$ with $\sigma^{2}=\frac{2}{N}\left(\frac{1}{m^{(j)}}+\frac{1}{m^{(k)}}\right)$ and $B_{t}$ a standard Brownian motion, we see that for two clusters on a periodic domain $[0,1]$, the timescale for the two clusters to merge is of order $N$. of the one dimensional Brownian motion in a bounded domain that a unique cluster will form almost surely in finite time on a timescale of order $t_{\rm coalesence}\asymp N$.
@@ 4.2. The extended Markovian dynamics with mass exchange
We now extend the model of the Markovian dynamics of the clusters by allowing the clusters to exchange mass. Under 1.1, we propose the following model, which we derive in Section 4.5: Consider the regime $\gamma\ell\gg 1$ and $\ell\ll 1$. We assume that at time $t$, there are $K(t)$ clusters with centres $X^{(1)}<\dots<X^{(K(t))}$ and masses $m^{(1)},\dots,m^{(K(t))}\in(0,1)$ such that $\sum_{j=1}^{K(t)}m^{(j)}=1$. The positions and the masses of the clusters behave as follows:
@@
  • Cluster merging: If two clusters with centres $X^{(j)}$ and $X^{(k)}$ meet, they merge into a single cluster of mass $m^{(j)}+m^{(k)}$ and position given by (4.6), so that the system (4.7) starts anew with the clusters $X^{(j)}$ and $X^{(k)}$ being replaced by a single cluster with starting position $X^{(j,k)}$ and mass $m^{(j)}+m^{(k)}$.
@@ 4.3. Application to the mean-field equation and dynamic metastability
The behaviour of the mean-field PDE (1.6) after initial clustering in the regime $\gamma\ell\gg 1$ and $\ell\ll 1$ is governed by the limit $N\to\infty$ of (4.7), so that the cluster centres are frozen in space and the mass exchange ODE system (4.7b) determines the dynamics. This is also reflected by the coalescence timescale $t_{\rm coalesence}\asymp N$. This suggests that after initial clustering, coarsening in the PDE (1.6) occurs by mass exchange between clusters which is described by (4.7b). Indeed, if we start the PDE with a multi-clustered initial state, then the ODE (4.7b) approximately describes the PDE evolution, as shown in Fig. 4.1 for the potential (1.4). Fig. 4.2 shows a similar comparison for the truncated piecewise parabolic potential with parameters $\alpha,\beta>0$, $a\in(0,1)$ given by
@@
For $\gamma\ell\gg 1$, the timescales $e^{\gamma\ell\Delta m}$ (see (LABEL:eq:clusters-masses:Eyring-Kramers) below) for a particle to leave a cluster are exponentially large in the cluster masses $m^{(j)}$, so that we expect dynamical metastability for the cluster mass dynamics. Indeed, assume, for instance, that we start the PDE with two clusters of masses $m_{a},m_{b}\approx\frac{1}{2}$, with $m_{a}>m_{b}$. In the regime $\gamma\ell\gg 1$, $\ell\ll 1$, we find from (4.7b) that mass is exchanged at a rate that is exponentially small in $\gamma\ell m_{a}$ and $\gamma\ell m_{b}$. Therefore, it will take exponentially long for the stronger cluster to absorb the smaller one. Indeed, we observe both for the PDE and the mass exchange ODE (4.7b) that the masses change only very little over a large amount of time to then suddenly collapse to a state in which only the initially heavier cluster remains, see Fig. 4.1.
@@
Observe that the model is only a valid approximate description of the PDE with a multi-clustered initial state if $X^{(j+1)}-X^{(j)}\gg\ell s_{w}$, that is, the initial centres of the clusters are sufficiently far apart, since otherwise deterministic merging phenomena may occur.
Moreover, the model (4.7b) is in general not a good approximation of the dissolution of clusters. Once one of the clusters has a mass smaller than the critical mass $m_{\rm crit}$ (see Section 3.6), then the dynamics of the particles of this cluster are not described by the Ornstein–Uhlenbeck approximation from [GPY17] and the cluster shape is in general not well approximated by (3.61). Moreover, in this situation, not only the behaviour of $w$ around zero but its general shape determines the behaviour and the exact evolution of mass. In particular, the description of the ODE (4.7b) is in general only valid until the first dissolution of a cluster.
@@
As before, we assume that for some fixed time $t\geq 0$, the particles (1.1) are arranged in clusters of masses $m^{(1)},\dots,m^{(K(t))}\in(0,1)$ and centres $X^{(1)}<\dots<X^{(K(t))}$. For simplicity, let us assume that the particles move on the real line $\mathbb{R}$ instead of the torus. Each cluster $X^{(j)}$ consists of approximately $m^{(j)}N$ particles. We would like to compute the timescale of a particle $X^{i}_{t}$ in the $j$-th cluster to leave the cluster and join another one, where $i\in\{1,\dots,N\}$ and $j\in\{1,\dots,K(t)\}$. The empirical measure $\mu^{N}_{t}:=\frac{1}{N}\sum_{i=1}^{N}\delta_{X^{i}_{t}}$ can be approximated by a linear combination of Gaussian densities centered at the cluster centres $X^{(k)}_{t}$ with variances $\sigma^{2}_{k}=\frac{\ell}{\gamma w^{\prime\prime}(0)m^{(k)}}$ and weights $m^{(k)}$:
@@
where $g(x;\mu,\sigma^{2})$ is the density of a normal distribution with mean $\mu$ and variance $\sigma^{2}$. We are interested in the event that the particle $X^{i}_{t}$ of cluster $j$ leaves this cluster and joins another cluster which can be assumed to be one of the neighboring clusters j1j-1 or $j+1$, since the event that a particle jumps to a cluster that is not a direct neighbor of the cluster where the particle has a probability of lower order. The particle $X^{i}_{t}$ moves in the effective potential
$\left(W_{\gamma,\ell}\ast\mu^{N}_{t}\right)(x)\approx V_{\rm eff}(x):=\sum_{k=1}^{K(t)}m^{(k)}\left(\gamma\ell w\left(\frac{{\color[rgb]{.75,.75,.75}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.75,.75}\pgfsys@color@gray@stroke{.75}\pgfsys@color@gray@fill{.75}{\bullet}}}{\ell}\right)\ast g({\color[rgb]{.75,.75,.75}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.75,.75}\pgfsys@color@gray@stroke{.75}\pgfsys@color@gray@fill{.75}{\bullet}};X^{(k)},\sigma_{k}^{2})\right)(x),$ (4.14)
that is, $X^{i}_{t}$ approximately follows the SDE ${\rm d}X^{i}_{t}=-V^{\prime}_{\rm eff}(X^{i}_{t}){\rm d}t+\sqrt{2}{\rm d}B^{i}_{t}$.
In order to approximate $V_{\rm eff}$ around its local minimum $X^{(j)}$, we replace $w(x)$ in (4.14) by the quadratic approximation $w(x)\approx-\Delta+w^{\prime\prime}(0)\frac{x^{2}}{2}$ for small $\lvert x\rvert$, where $\Delta:=-\inf w=-w(0)$. This is justified by the fact that the $\sigma_{k}^{2}$ is of order $\ell/\gamma$, while the interaction range of the potential $W_{\gamma,\ell}$ is of order $\ell\gg\sqrt{\ell/\gamma}$. Using this approximation, we obtain then
@@
for $x$ close to $X^{(k)}$, see also Fig. 4.3. Fix some $a\in\mathbb{R}$ inside the potential well of $V_{\rm eff}$ around $X^{(j-1)}$ which is close enough to the right boundary of this potential well and some $b\in\mathbb{R}$ inside the well of $V_{\rm eff}$ around $X^{(j+1)}$ which is close enough to the left boundary of this well. Let $\tau_{a,b}$ be the time of $X^{i}_{t}$ to hit $(-\infty,a]\cup[b,\infty)$ and set $u(x)=\mathbb{E}\left[\tau_{a,b}\nonscript\>\middle|\nonscript\>\mathopen{}X^{i}_{0}=x\right]$. By classical results [DZ01, Theorem 5.7.3] [Pav14, Sec. 7.3], $u$ satisfies the boundary value problem
@@
Figure 4.3. The effective potential $V_{\rm eff}$ for the Hegselmann–Krause prototype (1.4). The solid curve shows the convolution of $W_{\gamma,\ell}$ with a Gaussian mixture, the dashed curve shows the locally quadratic approximation (4.5) around the minima $X^{(j)}$ which is truncated by zero from above.
We also need the probability that a particle $X^{i}_{t}$ which starts at $x$ hits $a$ before hitting $b$ which is given by the equilibrium potential $h_{a,b}(x):=\mathbb{P}\left[\tau_{a}<\tau_{b}\nonscript\>\middle|\nonscript\>\mathopen{}X^{i}_{0}=x\right]$. Then, $h_{a,b}$ satisfies the boundary value problem given by (see e.g. [Ber13])